OpenStudy (anonymous):

(s-3 over 5)^2) i need step by step work done

jimthompson5910 (jim_thompson5910):

\[\Large \left(s-\frac{3}{5}\right)^2\] \[\Large \left(s-\frac{3}{5}\right)\left(s-\frac{3}{5}\right)\] \[\Large s*s+s\left(-\frac{3}{5}\right)+\left(-\frac{3}{5}\right)s+\left(-\frac{3}{5}\right)\left(-\frac{3}{5}\right)\] \[\Large s^2-\frac{3}{5}s-\frac{3}{5}s+\frac{9}{25}\] \[\Large s^2-\frac{6}{5}s+\frac{9}{25}\] So \[\Large \left(s-\frac{3}{5}\right)^2=s^2-\frac{6}{5}s+\frac{9}{25}\] for all values of 's'